Hello, I desperately need help with this problem. This problem is not from a college text but was made up by my math teacher. I'm not handy with a scientific calculator and have not been able to figure out how to get negative exponents to work equation. The problem looks like this:
You have decided to buy a used car. The cars from 4 dealers appear to be similar in price; however, the dealerships are offering different loan packages and you need to sit down & compare the financial advantages and disadvantages of each.
Using compount interest formula:
A=R[1-(1+i/n)raised to -nt power]
_____________________________
[ i/n ]
Where A stands for amount that you have to borrow from dealer, R represents the amount you will have to pay back per time compounded per year, i represents the annual % rate that you pay in interest, n represents the # of times per year that the interest would be compounded, and t represents the number of years that you would be making payment on the car.
Suppose you want to have the loan paid off in 4 years. How much would your car payment be in each case? Assume that the number of car payments that you make per year equals the number of interest periods per year. Details regarding dealers are below:
Amount Borrowed APR% # of Times Compounded Per Year
Dealer 1 $10,000 8.3% 12
Dealer 2 $9,500 8.6% 4
Dealer 3 $9,000 9% 6
Dealer 4 $10,500 8.2% 12
If I can get assistance with working the first dealer, I believe I could complete the calculations for dealers 2, 3, & 4. Thank you





Answer by stanbon(48551)

A=R[1-(1+i/n)raised to -nt power]/(i/n)
10000=R(1+0.083/12)^(-48)]/(0.0830/12))
10000(0.083/12)=R(1.00691667)^(-48)
830=R[0.71830781]
R=$1155.49
Cheers,
stan H.